OpenStudy (anonymous):

Simplify x/7x+x^2 Please help! thanks! I will give medals!

jimthompson5910 (jim_thompson5910):

hint: 7x+x^2 factors to x*(7+x)

OpenStudy (anonymous):

\[\frac{x }{7}x+x^2\text{=}\frac{x^2}{7}+x^2\text{=}\frac{8 x^2}{7} \]